OpenStudy (anonymous):

Hi, I was wondering if anyone could just be my tutor for just a little bot and show me how to make examples and solve for these: Product Property: log(A*B) = log(A) + log(B) Quotient Property: log(A/B) = log(A) - log(B) Power Property: log(A^b) = b*log(A)

OpenStudy (cwrw238):

ok try this one change to a single logarithm' log x + log y

OpenStudy (cwrw238):

use one of the laws of logs that you've written above to change this to log(something)

OpenStudy (cwrw238):

use the first one

OpenStudy (cwrw238):

yes or you could write it as log(xy) the multiply sign can be left out

OpenStudy (cwrw238):

now how about log(y) - log(x) write this as a single log

OpenStudy (cwrw238):

yes for the first part - thats log 42 of course and u r correct with the second part

OpenStudy (cwrw238):

now try this one write log (x^2) in another form

OpenStudy (cwrw238):

this uses your third law

OpenStudy (cwrw238):

yes - again no need to put the '*' 2 log(x) will do

OpenStudy (cwrw238):

now try one a little nore involved write the following as a single log log a + log b + 2 log c

OpenStudy (cwrw238):

thats ok

OpenStudy (cwrw238):

hint with this one: change the last log into log (something)

OpenStudy (cwrw238):

2 log c = log (c^2)

OpenStudy (cwrw238):

then do the final part

OpenStudy (cwrw238):

no it will be log (abc^2)

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

2 log (abc) = log (abc)^2

OpenStudy (cwrw238):

just go through the process indicating which property you are using

OpenStudy (cwrw238):

like the first one Convert to a single logarithm:- log x + log y By the Product Property log (a*b) = log a + log b (in reverse) log x + log y = log (x*y) or log (xy)

OpenStudy (cwrw238):

ok gotta go hope this has helped